OpenStudy (anonymous):

MEDAL AND FAN! Choose the equivalent system of linear equations that will produce the same solution as the one given below. 4x-2y=6 2x+y=5

OpenStudy (anonymous):

A) -4x-2y=10 -4y=4 B) -4x-5y=-1 -7y=5 C) 3x+2y=6 7x=12 D) 4x+2y=10 8x=16

OpenStudy (anonymous):

@Hotchellerae21 could you please help me?

OpenStudy (anonymous):

@hartnn could you please help me?

OpenStudy (anonymous):

4x + 2y = 10 8x = 16

OpenStudy (anonymous):

How did you do that?